# The scalar-torsion gravity corrections in the first-order inflationary models

### v1

@@ VI Inflationary models with linear dependence $r\sim(1-n_{S})$
Since, the zeroth order term in this expansion $r(0)=\beta_{0}=0$ from condition $r(n_{S}=1)=0$ corresponding to the flat Harrison-Zel’dovich spectrum [48, 49], we can rewrite expression (VI) in the following form
@@
The first term in expression (VI) makes the main contribution to the value of the tensor-scalar ratio. Thus, we can consider the dependence $r=r(1-n_{S})$ with a reasonable degree of accuracy at the first order
@@
to analyze the influence of non-minimal coupling on the parameters of inflationary models.
To obtain this dependence, we consider the inflationary models with minimal coupling and constant relation between slow-roll parameters during inflation
$\frac{\tilde{\delta}}{\tilde{\epsilon}}=\frac{\tilde{\delta}_{\ast}}{\tilde{\epsilon}_{\ast}}=const.$ (81)
@@
VI.1 Inflationary models with power-law potential and minimal coupling
@@
Further after substituting (91) into (VI.1), we obtain
@@
Thus from (96), we obtain the value of the scalar field at the crossing of the Hubble radius
@@
Also, from (92)–(93) and (97) we can write the slow-roll parameters at the crossing of the Hubble radius as
@@
From (71)–(72) and (98) we obtain the following expressions for the spectral tilt of the scalar perturbations and tensor-to-scalar ratio
@@
Thus, from equations (99)–(100) we obtain following dependencies of the tensor-to-scalar ratio from the spectral index of the scalar perturbations
@@
VI.2 Inflationary models with power-law potential and non-minimal coupling
Now, we estimate the influence of non-minimal coupling between the scalar field and torsion on the cosmological parameters. From equations (40)–(41) and (91) we obtain the following expressions for the potential and non-minimal coupling function
@@ VI.2 Inflationary models with power-law potential and non-minimal coupling
After substituting (105) and (107) into (88)–(90) we obtain
@@
Also, from (66)–(67) and (109) we get following dependence tensor-to-scalar ratio from spectral tilt of the scalar perturbations
@@
As one can see, for the case $n=0$ this dependence is reduced to (101).
Further, taking into account the observational constraint $r<0.032$ and restrictions (103)–(104) for the same observational values of spectral tilt $\tilde{n}_{S}=n_{S}=0.9663\pm 0.0041$, from (110) we obtain following constraints on the parametrization constant
$0.35<n<1,$ (111)
corresponding to the verifiable inflationary models with potential (105).
Also, we get the following possible values of the velocity of the scalar perturbations $0.70<c_{S}<$ 0.7 from expression (56). Thus, in the sub-luminal regime of the propagation of scalar perturbations, this model corresponds to the observational constraints (43)–(44). Therefore, for the case of scalar-torsion gravity, inflationary models with power-law potential can be verified by observational constraints on the values of the parameters of cosmological perturbations.
VI.3 Inflationary models with exponential potential
$\tilde{V}(\phi)=\tilde{V}_{0}\exp(-\alpha\phi),$ (112)
where $\tilde{V}_{0}$ and $\alpha$ are a some positive constants.
From (83)–(84) we obtain
$\tilde{\epsilon}=\tilde{\epsilon}_{\ast}=\frac{\alpha^{2}}{2}>0,$ (113)
Thus, from (71)–(72) and (114) we obtain
$\tilde{r}=8(1-\tilde{n}_{S}).$ (115)
After substituting $\tilde{n}_{S}\simeq 0.97$ into (115) one has $\tilde{r}\simeq 0.24$. Thus, the inflationary model with exponential potential (112) doesn’t satisfy the observational constraint (44) for the case of minimal coupling. For the case of non-minimal coupling, from equations (40)–(41) and (112) we get
$F(\phi)\simeq F_{0}\exp\left(-\frac{n\phi}{2(1+n)}\right),$ (117)
@@ VI.3 Inflationary models with exponential potential
After substituting (116) into (88)–(90) we obtain
that coincides with relation (75).
Further, from (66)–(67) and (120) we obtain
$r=8\sqrt{\frac{1-n}{1+n}}\,(1-n_{S}).$ (121)
Thus, for $n_{S}\simeq 0.97$ and $r<0.032$, from (121) we get $0.9650<n<$ 0.965. Also, for $0.9650<n<$ 0.965 corresponding velocity of the scalar perturbations (56) takes the values $0.130<c_{S}<$ 0.13.
Therefore, this inflationary model is verified by observational constraints only for the case of a very strong influence of the non-minimal coupling on the velocity of propagation of the scalar perturbations. For this reason, the inflationary models with power-law potential seem more realistic.
VII Conclusion
@@
We testify the influence of the non-minimal coupling of the scalar field and torsion on the cosmological parameters based on the power-law relationship between the coupling function and the Hubble parameter $F=(H/\lambda)^{n}$. This parametrization implies the possibility of a model-independent assessment of the effect of non-minimal coupling on cosmological parameters. Within the framework of this approach, this influence was parameterized by one constant parameter $n$. Also, we found restrictions on the value of parametrization constant $-1<n<1$. The influence of the non-minimal coupling on cosmological dynamics, scalar field potential, and parameters of cosmological perturbations was assessed by this approach.
@@
It was also shown that cosmological inflationary models implying a linear relationship between the tensor-scalar ratio and the spectral index of scalar perturbations can be verified by observational constraints on the values of cosmological perturbation parameters due to the influence of the non-minimal coupling between the scalar field and torsion. It was also noted that such an influence leads to a sub-luminal regime of the propagation of scalar perturbations and a decrease in the value of the tensor-scalar ratio for the models under consideration.
@@
Further development of the proposed approach is associated with its generalization to the case of the violation of relations $H=h^{1/(1+n)}$ and $\delta/\epsilon=const$. This generalization leads to a wider class of cosmological models including models with a super-luminal velocity of propagation of scalar perturbations and increasing tensor-to-scalar ratio.
@@ Acknowledgements
BM acknowledges Bauman Moscow State Technical University(BMSTU), Moscow for providing the Visiting Professor position during which part of this work has been accomplished.

### v3

@@ VII Inflationary models with linear dependence $r\sim(1-n_{S})$
In zeroth order of this expansion $r(0)=\beta_{0}=0$ we consider condition $r(n_{S}=1)=0$ for the flat Harrison-Zeldovich spectrum [84, 85]. Thus, we can rewrite expression (VII) in the following form
@@
The first term in expression (VII) makes the main contribution to the value of the tensor-scalar ratio. Thus, we can consider dependence
@@
as the condition of the first order inflationary models with negligible high-order terms in (VII).
The inflationary models of the second-order with dependence
$r\simeq\beta_{2}(1-n_{S})^{2},$ (108)
were considered, for example, in [57, 58, 59]. In this case $\beta_{1}=0$, and terms above the second order are negligible ones.
Now, we consider the inflationary models with minimal coupling and constant relation between slow-roll parameters during inflation
$\frac{\tilde{\delta}}{\tilde{\epsilon}}=\frac{\tilde{\delta}_{\ast}}{\tilde{\epsilon}_{\ast}}=constant,$ (109)
which follows from (91).
@@
VII.1 Inflationary models with exponential potential
$\tilde{V}(\phi)=\tilde{V}_{0}\exp(-\alpha\phi),$ (119)
where $\tilde{V}_{0}$ and $\alpha$ are a some positive constants.
From Eqs. (111)–(112), we obtain
$\tilde{\epsilon}=\tilde{\epsilon}_{\ast}=\frac{\alpha^{2}}{2}>0,$ (120)
Thus, from Eqs. (82)–(83) and Eq. (121) we obtain
$\tilde{r}=8(1-\tilde{n}_{S}).$ (122)
After substituting $\tilde{n}_{S}\simeq 0.97$ into Eq. (122) one can have $\tilde{r}\simeq 0.24$. Thus, the inflationary model with exponential potential (119) doesn’t satisfy the observational constraint Eq. (51) for the case of minimal coupling.
For the case of non-minimal coupling, from equations Eqs. (47)–(48) and Eq. (119) we get
$F(\phi)\simeq F_{0}\exp\left(-\frac{n\phi}{2(1+n)}\right),$ (124)
where
After substituting (123) into (116)–(118) we obtain
that coincides with relation (90).
Further, from (79)–(79) and (127) we obtain
$r=8\sqrt{\frac{1-n}{1+n}}\,(1-n_{S}).$ (128)
For $n_{S}\simeq 0.97$ and $r<0.032$, from (128) we get $0.9650<n<$ 0.965. Also, for $0.9650<n<$ 0.965 corresponding velocity of the scalar perturbations (63) takes the values $0.130<c_{S}<$ 0.13.
On the other hand, from conditions (104) and (121) one has $n=-2$. Thus, this model does not imply an exit from inflation and cannot be considered as a relevant one.
VII.2 Inflationary models with power-law potential and minimal coupling
@@
Further after substituting (129) into (VII.2), we obtain
@@
Thus from Eq. (134), at Hubble radius crossing, we obtain the value of the scalar field a
@@
From Eqs. (130)–(131) and Eq. (135) also, at the Hubble radius, the slow-roll parameters can be obtained as,
@@
From Eqs. (82)–(83) and Eq. (136) we obtain the following expressions for the spectral tilt of the scalar perturbations and tensor-to-scalar ratio
@@
Thus, from equations Eqs. (137)–(138) we obtain following dependencies of the tensor-to-scalar ratio from the spectral index of the scalar perturbations
@@
VII.3 Inflationary models with power-law potential and non-minimal coupling
Now, we estimate the influence of non-minimal coupling between the scalar field and torsion on the cosmological parameters. From Eqs. (47)–(48) and Eq. (129) we obtain the following expressions for the potential and non-minimal coupling function
@@ VII.3 Inflationary models with power-law potential and non-minimal coupling
After substituting Eq.(143) and Eq. (145) into Eqs. (116)–(118) we obtain
@@
Also, from Eqs. (79)–(79) and Eq. (147) we get following dependence tensor-to-scalar ratio from spectral tilt of the scalar perturbations
@@
For $n=0$ this dependence is reduced to Eq. (139).
Taking into account condition of exit from inflation (104) and expression (131), we obtain
$p=\frac{4(1+n)}{(2+n)}.$ (149)
After substituting (149) into (143)–(146) we obtain
$V(\phi)\simeq V_{0}\phi^{2},$ (150)
$F(\phi)\simeq F_{0}\phi^{\frac{2n}{2+n}}.$ (151)
Thus, under condition of exit from inflation one has quadratic potential (150) for any value of the constant $n$, which defines the influence of the non-minimal coupling between the scalar field and torsion.
Also, from (148)–(149) we get
$r=16\sqrt{\frac{1-n}{1+n}}\left(\frac{1-n}{4-n}\right)(1-n_{S}).$ (152)
Further, taking into account the observational constraint $r<0.032$ and restrictions Eqs. (141)–(142) for the same observational values of spectral tilt $\tilde{n}_{S}=n_{S}=0.9663\pm 0.0041$. From Eq. (148), we obtain the following constraints on the parametrization constant
$0.56<n<1,$ (153)
corresponding to the verifiable inflationary models with potential Eq. (143).
Thus, from (46) and (153) we get the following estimation of the constant parameter $\lambda\simeq 1/4$ in dependence (18).
Therefore, from (18) we obtain
$F(\phi(t))\simeq\left(4H(t)\right)^{n}.$ (154)
Also, we get the following possible values of the velocity of the scalar perturbations $0.530<c_{S}<$ 0.53 from expression (63). Thus, in the sub-luminal regime of the propagation of scalar perturbations, chaotic inflation with quadratic potential corresponds to the observational constraints (50)–(51) and exit from inflation (104) as well.
Therefore, the super-luminal regime of propagation of scalar perturbations cannot be realized in this cosmological model.
VII.4 Correspondence between exact and approximate cosmological solutions
The exact solutions of cosmological dynamic equations (18)–(21) for Hubble parameter (101) are
$\phi(t)=\frac{2}{2+n}\sqrt{\frac{2}{\mu}(1+n)}(\lambda-\mu t)^{1+\frac{n}{2}},$ (155)
$V(\phi)=V_{0}\phi^{\frac{2n}{2+n}}\left(\phi^{\frac{4}{2+n}}-B\right),$ (156)
$F(\phi)=\frac{V_{0}}{3}\phi^{\frac{2n}{2+n}},$ (157)
@@ VII.4 Correspondence between exact and approximate cosmological solutions
$V_{0}=\frac{3\mu(2+n)^{2}}{8\lambda^{n}(1+n)},$ (158)
$B=\frac{\mu}{3}(1+n)\left(\frac{8(1+n)}{\mu(2+n)^{2}}\right)^{\frac{2}{2+n}}.$ (159)
As one can see, under condition $B\ll\phi^{\frac{4}{2+n}}$ exact form of potential (158) corresponds to (150) for the slow-roll approximation.
From (103) we obtain following relation
$H=\left(\frac{\mu}{\epsilon}\right)^{1/2},$ (160)
and at the crossing of the Hubble radius
$H_{\ast}=\left(\frac{\mu}{\epsilon_{\ast}}\right)^{1/2}.$ (161)
Based on (160), we can rewrite the expression for the scalar field (155) as follows
$\phi(\epsilon)=\frac{2}{2+n}\sqrt{\frac{2}{\mu}(1+n)}\left(\frac{\mu}{\epsilon}\right)^{\frac{2+n}{4}}.$ (162)
From (153), (159) and (162) we have
$\frac{\phi^{\frac{4}{2+n}}}{B}=\frac{C}{\epsilon}\sim\epsilon^{-1},$ (163)
where $1.91.5<C<$ 1.9.
Thus, condition $B\ll\phi^{\frac{4}{2+n}}$ is actual while slow-roll condition $\epsilon\ll 1$ is satisfied.
Therefore, we can consider potential (156) as
$V(\phi)\sim\phi^{2},$ (164)
during slow-roll regime.
At the end of inflation with $\epsilon=1$ the second term in potential (156) becomes significant.
Also, from expression (79) taking into account $\delta_{\ast}=0$, we obtain
$\epsilon_{\ast}=\frac{1-n_{S}}{4-n}.$ (165)
From (153) and (165) for $n_{S}\simeq 0.97$ we obtain the following values of the first slow-roll parameter at the crossing of the Hubble radius $0.010.0087<\epsilon_{\ast}<$ 0.01. Therefore, at the crossing of the Hubble radius slow-roll conditions are satisfied.
From definitions of e-folds number implying $\dot{N}=H$ and the first slow-roll parameter (39) for Hubble parameter (101) we get
$\frac{dN}{d\epsilon}=\frac{dN}{dt}\frac{dt}{d\epsilon}=\frac{H}{\dot{\epsilon}}=\frac{1}{2\epsilon^{2}}.$ (166)
Thus, the e-folds number between the crossing of the Hubble radius and the end of inflation is
$\tilde{N}=\frac{1}{2}\int^{\epsilon_{e}=1}_{\epsilon_{\ast}}\frac{d\epsilon}{\epsilon^{2}}=\frac{1}{2}\left(\frac{1}{\epsilon_{\ast}}-1\right)$
$\simeq\frac{1}{2\epsilon_{\ast}}=\frac{4-n}{2(1-n_{S})}.$ (167)
From (153) and (VII.4) for $n_{S}\simeq 0.97$ we obtain the following values of the e-folds number between the crossing of the Hubble radius and the end of inflation
$57\displaystyle 50<\tilde{N}<$ 57 . (168)
Also, from (79)–(79) we obtain the following expressions for the spectral index of the scalar perturbations and the tensor-to-scalar ratio
$1-n_{S}=\frac{4-n}{2\tilde{N}},$ (169)
$r=\frac{8(1-n)}{\tilde{N}}\sqrt{\frac{1-n}{1+n}},$ (170)
which leads to the same dependence (152).
From (79) and (165) we get
${\mathcal{P}}_{S}=\frac{(2-n)(4-n)H^{2-n}_{\ast}}{8\pi^{2}(1-n^{2})\sqrt{\frac{1-n}{1+n}}\,(1-n_{S})}.$ (171)
Thus, from (153) and (171) for ${\mathcal{P}}_{S}=2.1\times 10^{-9}$ and $n_{S}\simeq 0.97$ we obtain following scale of inflation
$0<H_{\ast}<3\times 10^{-7},$ (172)
in the chosen system of units.
Also, from (154) and (172) we obtain the following constraint on the values of the coupling function at the crossing of the Hubble radius
$0<F_{\ast}<5\times 10^{-4}.$ (173)
From (153), (161), (165) and (172) we obtain
$0<\mu<8\times 10^{-16}.$ (174)
From (68)–(70), (153) and (158)–(159) we get
$V_{0}<5.4\times 10^{-15},$ (175)
$B<4.2\times 10^{-4}.$ (176)
Also, from (III), (173) and (174) the value of the potential of the scalar field at the crossing of the Hubble radius is
$V_{\ast}=F_{\ast}\left(3H^{2}_{\ast}-\mu(1+n)\right)<1.3\times 10^{-16}.$ (177)
Kinetic energy of the scalar field (23) at the crossing of the Hubble radius is
$X_{\ast}=F_{\ast}\mu(1+n)<8\times 10^{-19}.$ (178)
Thus, the slow-roll condition $V_{\ast}\gg X_{\ast}$ is satisfied in this inflationary model.
In natural system of units [1] potential (177) can be estimated as follows
$V_{\ast}<1.36\times 10^{-16}\times M^{4}_{P}=4.3\times 10^{57}~{}(GeV)^{4},$ (179)
where reduced Planck mass is $M_{P}=\left(8\pi G\right)^{-1/2}=2.4\times 10^{18}$ GeV.
Corresponding energy scale of inflation is
$V^{1/4}_{\ast}<2.6\times 10^{14}~{}GeV<\Lambda_{GUT}<M_{P},$ (180)
where $\Lambda_{GUT}\sim 10^{16}$ GeV is GUT scale which is considered as the upper bound on the energy scale of inflationary models [1].
Also, from (153), (162), (165) and (174) we have following estimates for the value of a scalar field
$0<\phi_{\ast}<0.22,$ (181)
$0.01\displaystyle 0<\phi_{e}<$ 0.01 . (182)
Now, we will check to what extent this model satisfies Swampland criteria [91, 93, 94]
$\left|\Delta\phi\right|=\left|\phi_{e}-\phi_{\ast}\right|\leq{\mathcal{O}}\left(1\right),$ (183)
$\left|\frac{V_{,\phi}}{V}\right|_{\phi=\phi_{\ast}}\geq{\mathcal{O}}\left(1\right).$ (184)
Satisfaction of the Swampland criteria means that the cosmological model can be regarded as a low-energy effective one in quantum gravity [91, 93, 94].
On the basis of expressions (164) and (181)–(182) we obtain
$\left|\Delta\phi\right|<0.21<{\mathcal{O}}\left(1\right),$ (185)
$\left|\frac{V_{,\phi}}{V}\right|_{\phi=\phi_{\ast}}=\frac{1}{\phi_{\ast}}\simeq 4.5>{\mathcal{O}}\left(1\right).$ (186)
Thus, the proposed model of cosmological inflation based on the scalar-torsion gravity with potential (156) and coupling function (157) satisfies the Swampland criteria, observational constraints on the parameters of cosmological perturbations and the condition of the graceful exit from inflation.
VIII Conclusion
@@
We testify the influence of non-minimal coupling between scalar field and torsion on the cosmological parameters based on power-law relationship between coupling function and Hubble parameter $F=(H/\lambda)^{n}$. The novelty of the proposed approach is the possibility of a model-independent assessment of non-minimal coupling effect on cosmological parameters. The non-minimal coupling influence was parameterized by one constant parameter $n$. Also, we found restrictions on the value of parametrization constant $-1<n<1$. The influence of non-minimal coupling on cosmological dynamics, scalar field potential, and cosmological perturbations parameters was assessed by this approach.
@@
We have shown that inflationary models implying a linear relationship between the tensor-scalar ratio and spectral index of scalar perturbations can be verified by observational constraints on the values of cosmological perturbation parameters by influence of the non-minimal coupling between scalar field and torsion without the need to take into account the Galilean field self-interaction. These influence leads to a sub-luminal regime of the propagation of scalar perturbations and a decrease in the value of the tensor-scalar ratio for inflationary models under consideration. A non-trivial property of the power-law relationship $F=(H/\lambda)^{n}$ is the constant velocity of propagation of scalar perturbations in inflationary models under consideration. Also, we note that expressions (18) and (27) define the relations of the constant parameters $n$ and $\lambda$ with $c_{S}$. Thus, in this case, the parametrization of the influence of non-minimal coupling between the scalar field and torsion can be performed only based on the propagation velocity of scalar perturbations $c_{S}=constant$ following expressions (68)–(70).
@@
Further development of the proposed approach is associated with its generalization to the case of the violation of relations $H=h^{1/(1+n)}$ and $\delta/\epsilon=const$. This generalization leads to a wider class of cosmological models beyond the considered type of inflationary model.
Finally, we note that taking into account the Galilean field self-interaction can also significantly increase the number of the possible verifiable inflationary models.
@@ Acknowledgements
BM acknowledges Bauman Moscow State Technical University(BMSTU), Moscow for providing the Visiting Professor position during which part of this work has been accomplished. LKD acknowledges the financial support provided by University Grants Commission (UGC) through Senior Research Fellowship UGC Ref. No.: 191620180688 to carry out the research work.